Given the contingency table shown here, find PIV or ). Vehicle Type Car (C) Minivan (M) Full-Size Van (F) SUV (V) Truck (T) Cfind P(V or S):

answer: a. .4300 b. .0475 c. .5100 d. .4775
Answer #1

P(V) = No. of SUV/Grand Total = 72/300

P(S) = Total vehicles in Somerset/ Grand Total = 100/300

P(V and S) = P(V\cap S) = No of SUVs in Somerset/300 = 19/300

We know,

P(V or S) = P(V\cup S) = P(V) + P(S) - P(V\cap S)

=\frac{72}{300} +\frac{100}{300}-\frac{19}{300}=0.51

Ans: Option C